UT20090625 Telescope Performance Verification
with LUCIFER at LFBG

Observers: W. Seifert, N. Ageorges, V. Knierim, J. Hill (LBT)
Telescope Operator: A. Ceranski
Software Support: M. Delapena, (Tucson)
Instrument Support: T. Shih
Telescope Support: J. Little

Summary

Lost whole night to clouds.

Preparations

Dave A. did some servo tuning on the LFBG rotator.

Butch fixed the failed pop-off valve that was causing problems on the SX air dryer.

Jim D. repaired the disk on lbtmu304.

John H. toured 3 dozen astronomy campers around the facility.

Plenty of time for afternoon and evening flexure tests with LUCIFER.

Update LUCIFER Collimation Models

I updated the LUCI collimation models to compensate the Z7=+4000 nm coma correction that we needed at the end of last night when we were in decent equilibrium. This means moving M1 by dY=+3.414 and dRX=+36.68, but as we discovered last night this runs us into the +Y limit. So I also translated both the primary and secondary mirrors by -3.5 mm in dY. So the net global offsets are as follows:
SXPMLUCIFERCollimation.dat: 0.000 -0.100 0.000 36.680 0.000 0.000 Global Offsets
SXSMLUCIFERCollimation.dat: -7.010 -3.879 0.000 148.650 181.800 0.000 Global Offsets

What I don't understand is why the coma correction has changed so much since Dave measured the collimation model with IRTC on May 1 when the temperature was 7 degC. Those corrections were apparently still good for IRTC 2 weeks ago. Does LUCIFER bend the gallery enough to change the direction of the rotator axis in a significant way? (I thought we had ruled that out in past measurements.) This is a good thing to remeasure in the September startup time.

I've resisted it so far because it makes the collimation corrections harder to keep track of, but I'm thinking that we probably need to introduce a Y temperature term that translates M1 and M2 together between summer and winter.

Details

Clouds and monsoon showers early. Waiting for it to clear off.

Closed-dome Preset Tests

06:37 pointto BS9170 ACTIVE PA=0 - failed for lack of counts

06:42 pointto BS9170 ACTIVE PA=90 -

06:44 pointto BS9177 ACTIVE PA=90

06:47 pointto BS9177 TRACK PA=90

07:20 back to zenith for more LUCIFER tests

still waiting..............still cloudy. It didn't actually rain until after sunrise.

Appendix

top - 07:46:24 up 21 days,  7:54,  3 users,  load average: 6.06, 5.09, 4.44
Tasks: 131 total,   1 running, 130 sleeping,   0 stopped,   0 zombie
Cpu(s):  0.3%us,  0.2%sy,  0.0%ni, 99.3%id,  0.1%wa,  0.0%hi,  0.1%si,  0.0%st
Mem:   8313508k total,  5849556k used,  2463952k free,   148612k buffers
Swap:  2031608k total,       96k used,  2031512k free,  5213788k cached

  PID USER      PR  NI  VIRT  RES  SHR S %CPU %MEM    TIME+  COMMAND                                                                   
26612 telescop  15   0  6456 2688 2452 S  1.0  0.0 152:00.19 gshmserver                                                                
 8703 telescop  15   0  318m  30m 6044 S  0.3  0.4  21:55.30 PCS                                                                       
26874 telescop  18   0  908m 142m 9.8m S  0.3  1.8  95:53.51 IIF                                                                       
    1 root      15   0  2064  620  536 S  0.0  0.0   0:01.27 init         

-- JohnHill - 25 Jun 2009
Topic revision: r5 - 26 Jun 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback